# Constructing equivalences between fusion categories of quantum groups and of vertex operator algebras via quantum gauge groups

### v2

@@ Abstract.
We summarize our results in [11] on a problem that Sergio Doplicher posed in the 90s in the setting of conformal nets, and on a related problem posed by Yi-Zhi Huang in the setting of vertex operator algebras, see e.g. [55]. The latter asks to find a direct construction of an equivalence due to Finkelberg between the quantum group modular fusion category at certain roots of unity and the fusion categories of affine vertex operator algebras at positive integer levels. The former asks for the construction of a quantum gauge group and field algebra for the category of localized endomorphisms of a conformal net, extending Doplicher-Roberts theory of the compact gauge group and field algebra in high dimensional algebraic QFT. In this summary we mostly focus on the affine VOA aspects (corresponding to the loop group conformal nets).
We solve Huang’s problem for all the Lie types for which it is known that the centralizer algebras of the non-negligible tensor powers of a given generating object $V$ for the quantum group fusion category (a list of $V$ for which fusion tensor powers are well defined has been given by Hans Wenzl in [91] for all the Lie types), is generated by the representation of the braid group associated to the $R$-matrix, and for Huang-Lepowsky ribbon braided tensor structure for the vertex operator algebras. In the classical type $A$, this property becomes the classical Schur-Weyl duality for the vector representation, and was used by Doplicher and Roberts to construct ${\rm SU}(d)$ as a special case of their unique compact gauge group and field algebra in algebraic QFT. Presently, known results in the literature on this generating property allow to construct our equivalence for the Lie types $A$, $B$, $C$, $D$, $G_{2}$. Our approach may be regarded in the setting of Connes noncommutative geometry.
Lack of knowledge of the braid group generation property (i.e. for the Lie types $E$ and $F$) allows us to show equivalence with Huang-Lepowsky ribbon braided tensor structure for all the ribbon morphisms, and part of those braiding morphisms, and associativity morphisms for which $V$ takes one or two components, respectively.
Our work also gives a unitary modular tensor category structure on the linear category of modules of the Zhu algebra associated to the minimum energy functor (or on the vertex operator algebra) for all the Lie types, equivalent to that from quantum groups. This algebra becomes a unitary weak quasi-Hopf algebra, solving a problem posed by Frenkel and Zhu for all the Lie types [35].
Our main tool is the construction of two semisimple quantum gauge groups and an equivalence of their representation categories: The quantum group fusion category provides a unitary coboundary weak Hopf algebra associated to Wenzl functor in [91]. We then apply a de-quantization continuous curve of Wenzl and a Drinfeld twist method to reach the Zhu algebra. The structures we find in both contexts extend Doplicher-Roberts symmetric functors for rigid symmetric tensor categories with simple unit which play a central role in their duality theory.
@@ 1. Introduction
On the side of CFT, our main interest in this paper is the Wess-Zumino-Witten model. Modular, or more generally ribbon rigid braided tensor category constructions involved many authors among physicists and mathematicians. A main difficulty in constructing tensor categories from models of conformal field theories is the construction of the associativity morphisms compatible with naturally assigned braiding morphisms via the two hexagonal equations.
Let ${\mathfrak{g}}$ be a complex simple Lie algebra. This problem was first solved by Drinfeld in [23], [25] [26] for the Drinfeld category with the introduction of the notion and construction of non-trivial examples of quasi-Hopf algebras over $U({\mathfrak{g}})[[h]]$, via the Knizhnik-Zamolodchikov differential equation which is satisfied by the correlation functions of the WZW model. Drinfeld used this equation to construct the associativity morphisms compatible with naturally assigned braiding morphisms.
Furthermore Drinfeld established his remarkable Drinfeld-Kohno theorem expanding an earlier theorem by Kohno [62] stating that the representation of the braid group arising from monodromy of the KZ equation is equivalent to that arising from the action of the $R$-matrix of $U_{h}({\mathfrak{g}})$ with respect to the vector representation, for the classical Lie types. Drinfeld-Kohno theorem connects via a Drinfeld twist, the quasi-Hopf algebra $U({\mathfrak{g}})[[h]]$ with the quantized universal enveloping algebra $U_{h}({\mathfrak{g}})$ which is a Hopf algebra. In this case, rigidity of Drinfeld category is derived from the equivalence constructed in Drinfeld-Kohno theorem and rigidity of the representation category of $U_{h}({\mathfrak{g}})$.
Major work was carried out in a series of papers by Kazhdan and Lusztig in the mid 90s in the setting of affine Lie algebras. The authors construct rigid braided tensor structures for certain module categories of affine Lie algebras associated to ${\mathfrak{g}}$ at levels excluding the positive rationals. In their work they consider both generic and non generic values of the levels. In particular they consider the negative integer levels $-k$, $k\in{\mathbb{N}}$, and proved a braided tensor equivalence of their category ${\mathcal{O}}_{-k}$ with the non-semisimple category of certain modules of the Drinfeld-Jimbo quantum group $U_{q}({\mathfrak{g}})$ at the same type of roots of unity $q=e^{i\pi/d\ell}$, with the exception of some exceptional cases at low levels [59]. The order $\ell$ appearing in the root of unity and the positive integer $k$ are related by a well known equality, see e.g. Def. 19.1 and 20.2 in [11] and references therein.
@@
The study of how unitarization from the quantum group fusion categories [91] was useful to derive unitarization in representation categories of vertex operator algebras via an explicit equivalence that passes through weak quasi-Hopf algebras was important for us since the early stages of applications of the theory of weak quasi-Hopf algebras [11], and further motivated us to study the above problem. Indeed one of our first results on unitarization was that weak quasi-Hopf algebras associated to fusion categories known to be equivalent allow to transport the unitary structure of one category to unitarize the other. In the process, knowledge of the explicit equivalence gives an understanding of the transported unitary structure. We have results on the structure of the unitary structure of the affine vertex operator algebra at positive integer levels transported by our equivalence, which asserts that it must correspond to the classical compact real form, and necessarily equals the unitary structure previously directly constructed by Gui, see [11] for more details and references.
@@
Rigidity, called charge conjugation in this setting, describes particle-antiparticle duality. In their papers rigidity is derived as a weak limit of inner automorphisms induced by transfer operators moving the support of the given localized endomorphism $\rho$ to infinity. This gives a left inverse $\Phi_{\rho}$ of $\rho$, which is a completely positive map. From this the conjugate localized endomorphism $\overline{\rho}$ is derived. The left inverse may be described in the category as a map induced by the adjoint action of a solution of the conjugate equations on a intertwiner translated to the right by the tensor product. It was shown that charge conjugation implies deep and useful properties for the classification of the statistics parameter (which is the scalar $\Phi_{\rho}(\varepsilon)$, with $\varepsilon$ the action of the generator of the permutation group in the algebra of intertwining operators of the tensor powers of $\rho$), and of charges into charges of para-Bose or para-Fermi type.
Because of the representation of the permutation group in the category, together with their results on the classification of the statistics parameter, which was shown to the the reciprocal of a positive integer $d$ (the order of the parastatistics) up to a sign (Bose or Fermi cases), and the $C^{*}$-structure, in the 70s it arose the question of whether that was the category of representations of a compact group, because those data arising from a single localised endomorphism were also satisfied by the representation category of a compact Lie subgroup of ${\rm SU}(d)$, that would be a gauge group of global symmetries. They considered the abelian case in the 70s [17]. In that case, every object had dimension $1$, and in modern terminology these categories are referred to as pointed in e.g. [27].
@@
Our approach may be regarded in the setting of Connes noncommutative geometry [13], and follows naturally our previous work in the setting of compact quantum groups, see also the introduction in [12], [14]. In [76], [77], [81], [78], [79], [80], [82] we have developed at the level of rigid tensor $C^{*}$-categories with simple unit an analogue of Doplicher-Roberts construction of a compact gauge group and field algebra, or more precisely the analogue of [21] in the setting of compact quantum groups. The analogy cannot be performed in a strict way due to new difficulties in the quantum case, see the mentioned papers and [11]. One important difference was that for rigid braided tensor categories we do not have a notion of braided tensor functor beyond the symmetric ones, because the category of vector spaces does not admit a braided symmetry. Giving up a compatibility property in the symmetric case implies lack of uniqueness of the functor, and thus of the gauge group. This was noticed early as an easy consequence from their construction.
@@
To treat the difficulty of finding a replacement of Doplicher-Roberts symmetric functors, in [10] and [11] we have taken into account Drinfeld notion of coboundary symmetry $\overline{R}$, which plays a role in the construction of Drinfeld twist in Drinfeld-Kohno theorem [25]. This is a symmetry of the category obtained modifying the braided symmetry defined by the $R$-matrix by a square root of the $2$-coboundary defined by the ribbon structure, and by construction the coboundary symmetry is an involutive symmetry, similarly to the braided symmetry. This symmetry finally allows for a generalization of the notion of symmetric functor. Moreover we have followed both the original choice by Mack and Schomerus [64] of a given dimension function corresponding to non-negligible tensor powers of the vector representation of quantum ${\mathfrak{s}l}_{2}$ at roots of unity and also the study of classification of compact quantum groups corresponding to a given representation category of deformation compact quantum groups by Neshveyev and Yamashita [74], [75]. They both considered a specific positive integer valued dimension function associated to the given category of the quantum group, which was the classical dimension function.
@@
This suggested to us to consider in [10] and [11] a specific dimension function on a functor related to classical data, and this was first Wenzl functor [91] for the quantum group fusion category now at roots of unity, and then the minimum energy functor for the WZW model [98] in the setting of vertex operator algebras. It was particularly important for us that in [91] Wenzl describes an explicit quantization curve associated to his functor that leads to an algebra that identifies with the Zhu algebra for that model in the setting of vertex operator algebras. Since in [10] for the type $A$ case we had constructed a natural quantum group structure on a semisimple associative algebra from the quantum group fusion category endowed with Wenzl functor, compatible with the tensor structure, this arose the question of constructing another natural quantum group structure on the Zhu algebra at the time of [10], following the quantization curve on the unitary structure of tensor products with the aim of making it classical. This involved in [11] the construction of a Drinfeld twist in our setting, the above mentioned notion that Drinfeld had introduced to study equivalence of the Drinfeld category with the quantum group category in a formal setting. We have done this for all the Lie types via our version of Drinfeld-Kohno Theorem 28.4 and applications to affine vertex operator algebras at positive integer level, via Zhu linear equivalence, in [11].
@@
We refer to Sect. 23 in [11] for an extension of certain results by Neshveyev and Yamashita in the setting of quantum group fusion categories of type $A$ based on Kazhdan-Wenzl theory at roots of unity in our case, and on the weak Hopf algebras constructed in [10], in place of the compact quantum groups. This section is important for us as it suggested that in general the associativity morphisms in a category arising from CFT may be of a simple nature from the viewpoint of the associated quantum gauge groups that need to be constructed, and this finally proved to be the case, we shall come back to this with more precision.
@@
On the other hand, in the process of extending Doplicher-Roberts theory in the setting of compact quantum groups, we had started by embedding in a tensorial way part of the morphisms of the category into the Hilbert spaces [76], [78], [82], and in doing this we had already made a choice of a dimension function, and from that choice we had constructed a tensor functor from that subcategory. To consider the full category, the analogue Doplicher-Roberts compact gauge group is described by an ergodic action of a larger quantum group on a quantum homogeneous space described by a $C^{*}$-algebra, rather than a quantum group itself, due to lack of commutativity of this $C^{*}$-algebra [78]. This property is likely to hold also in other models arising from conformal field theory starting from the minimum energy functor, unless one can show that certain known obstructions already met in [77] for compact quantum groups, vanish, we shall come back to this in the outlook section 13.
@@
However, for the WZW models arising from affine Lie algebras at positive integer levels in [10] for the type $A$ case, and in [11] for all the Lie types, we obtain genuine quantum groups as quantum gauge groups, described as semisimple weak Hopf algebras with $R$-matrix, $C^{*}$-structure explicitly associated to Drinfeld coboundary matrix, and with unitary representations of the braid group, derived from the quantum group fusion category. We prove that these weak Hopf algebras have representation category equivalent to both the starting fusion category and also to that of the corresponding affine vertex operator algebra for the classical Lie types and $G_{2}$. Constructing the equivalence from the former to the latter in the setting of semisimple weak quasi-Hopf algebras as we do in [11] is our proof of Kazhdan-Lusztig-Finkelberg theorem, see Theorem 2.1.
@@
Among our first references in the context of conformal field theory important for us to start to set up the connection with the work of Wenzl in quantum groups, is the work by Wassermann [90] in the setting of affine Lie algebras, where the key concept is played by the notion of primary field, and its initial term, which we use in a fundamental way to connect quantum group fusion category to representations of the Zhu algebra and their special tensor products with the fundamental representation, see Sect 33 in [11]. Most importantly, we describe a connection between Wenzl and Wassermann work via Wenzl quantisation curve in theorem 33.7, to which we refer the reader also for important historical references. The connection from affine Lie algebra setting to Huang-Lepowsky tensor product theory may be found in Sections 35 and 36.
@@
Our abstract duality methods need new further tools as we shall see from the theory of quantum groups, mainly from the work of Drinfeld and Wenzl, that are not available in the theory of groups. However, we shall find that in both contexts, quantum groups at roots of unity and affine vertex operator algebras at positive integer levels, both representation categories are described by functors extending the symmetric tensor functors in Doplicher-Roberts abstract duality theory, see Sect. 27 in [11]. We call these structures hermitian coboundary weak quasi-Hopf algebras, as the hermitian structure is given by Drinfeld coboundary matrix, which is a modification of the $R$-matrix with a square root of the ribbon structure.
@@
We notice on the other hand that a stronger property is needed to establish the connection, than the purely categorical characterisation of functors in Sect. 27. This property is expressed in terms of quantum groups, and the first example, in a non-semisimple setting, is the hermitian structure of $U_{q}({\mathfrak{g}})$ at a root of unity by Theorem 29.9. This property says that the -involution enjoys a precise anticommutation relation (rather than commutation relation as is familiar in Hopf $C^{*}$-algebras) with the coproduct and was first noticed in [91]. We axiomatize this property in section 28, as a compatibility property of hermitian coboundary weak quasi-Hopf algebras. Further examples derived from this examples are our weak Hopf and quasi-Hopf algebras, see Sections 29, 30, 31. This is not a direct consequence of the corresponding property of $U_{q}({\mathfrak{g}})$, as its coproduct is not compatible with the fusion tensor product.
@@
As above motivated, we shall derive more structure than rigidity and modularity, such as unitarization of the vertex operator algebra representation category from the corresponding quantum group fusion category, via the explicit equivalence, see Sects. 21, 22, 24 in [11]. To this aim it is important that our equivalence passes through the weak quasi-Hopf algebras, and the fact that these are naturally associated to the categories. Indeed it was known that a mere tensor equivalence between the categories would not suffice to transport the unitary structures see [43]. Sect. 15 in [11] treats this for general tensor categories that admit submultiplicative dimension functions (called weak dimension functions).
@@
Moreover our equivalence, which is given by an analogue of the Drinfeld twist in [23], [25], at the level of the associated weak Hopf or quasi-Hopf algebras gives an understanding of the unitary structures of representation categories of affine vertex operator algebras as classical compact real forms, or untwisted. We refer to the outlook section for some more information on possible extension of our viewpoint to models beyond the affine vertex operator algebras at positive integer levels, see also [11].
@@
Our main result is Theorem 2.1. In the subsequent sections we outline the plan of our proof. A main idea that we follow is to consider the fundamental representation $V$ of ${\mathfrak{g}}$ for each Lie type as in [91] where the author explains its generating property and good fusion structure with respect to tensor products with an arbitrary irreducible representation allowing a well-defined truncation procedure to obtain non-negligible objects, with the aim of studying unitarizability of the fusion category. It should also be noted that a main application of that paper is construction of inclusion of subfactors of finite Jones index for all Lie types, extending his previous work.
@@
For all the Lie types we obtain complete identification for the tensor product structure, all the ribbon structure, and those braiding and associativity morphisms where one or two variables are the fundamental representation $V$, respectively, with Huang-Lepowsky ribbon braided tensor structure. These are parts (a), (b) of Theorem 2.1.
@@
In [10], [11], and also in this paper, weak Hopf algebras are understood with a different meaning. While a different name would be helpful to avoid confusion, so far we have not been able to find a good name for our algebras to clarify the situation, and we use the term weak Hopf algebras in the sense defined in Sect 9 in [11].

### v4

@@ Abstract.
We explain the structure of proof of our results on a problem that Sergio Doplicher posed in the 90s in the setting of conformal nets, and on a related problem posed by Yi-Zhi Huang in the setting of vertex operator algebras. The latter asks to find a direct construction of an equivalence due to Finkelberg between the quantum group modular fusion category at certain roots of unity and the fusion categories of affine vertex operator algebras at positive integer levels. The former asks for the construction of a quantum gauge group and field algebra for the category of localized endomorphisms of a conformal net, extending Doplicher-Roberts theory of the compact gauge group and field algebra in high dimensional algebraic QFT. In this summary we mostly focus on the affine VOA.
We solve Huang’s problem for all the Lie types for which it is known that the centralizer algebras of the non-negligible tensor powers of a given generating object $V$ for the quantum group fusion category (a list of $V$ for which fusion tensor powers are well defined has been given by Hans Wenzl for all the Lie types), is generated by the representation of the braid group associated to the $R$-matrix, and for Huang-Lepowsky ribbon braided tensor structure for the vertex operator algebras. In the classical type $A$, this property becomes the classical Schur-Weyl duality for the vector representation, and was used by Doplicher and Roberts to construct ${\rm SU}(d)$ as a special case of their unique compact gauge group and field algebra in algebraic QFT. Presently, known results in the literature on this generating property allow to construct our equivalence for the Lie types $A$, $B$, $C$, $D$, $G_{2}$. Our approach may be regarded in the setting of Connes noncommutative geometry.
We also solve a problem by Frenkel and Zhu on the construction of a (unitary) weak quasi-Hopf algebra structure on the Zhu algebra, associated to the minimum energy functor for all the Lie types, equivalent to that from quantum groups.
Our main tool is the construction of a unitary coboundary weak Hopf $C^{*}$-algebra in a sense extending Drinfeld coboundary quasi-Hopf algebras on a formal variable, associated to Wenzl functor for the quantum group fusion category, and of a Drinfeld twist. We apply a de-quantization continuous curve of Wenzl and a Drinfeld twist method to reach the Zhu algebra. The two Drinfeld coboundary symmetries we find in both contexts extend Doplicher-Roberts symmetric functors for rigid symmetric tensor categories with simple unit which play a central role in their duality theory.
@@ 1. Introduction
On the side of CFT, our main interest in this paper is the Wess-Zumino-Witten model. Modular, or more generally ribbon rigid braided tensor category constructions involved many authors among physicists and mathematicians. A main difficulty in constructing tensor categories from models of conformal field theories is the construction of the associativity morphisms compatible with naturally assigned braiding morphisms via the two hexagonal equations. Another main difficulty is to show rigidity of the category.
Let ${\mathfrak{g}}$ be a complex simple Lie algebra. These problems were first solved by Drinfeld in [25], [27] [28] for the Drinfeld category with the introduction of the notion and construction of non-trivial examples of quasi-Hopf algebras over $U({\mathfrak{g}})[[h]]$, via the Knizhnik-Zamolodchikov differential equation which is satisfied by the correlation functions of the WZW model. Drinfeld used this equation to construct the associativity morphisms compatible with naturally assigned braiding morphisms. Furthermore he established rigidity of this category by showing a remarkable equivalence with the module category arising from the quantized universal enveloping algebra $U_{h}({\mathfrak{g}})$, which is a Hopf algebra. To construct the equivalence, he used Drinfeld coboundary matrix ${\overline{R}}$, which is a modification of the $R$-matrix by a square root of the coboundary defined by the ribbon structure.
This is his remarkable Drinfeld-Kohno theorem, expanding an earlier theorem by Kohno [69] stating that the representation of the braid group arising from monodromy of the KZ equation is equivalent to that arising from the action of the $R$-matrix of $U_{h}({\mathfrak{g}})$ with respect to the vector representation, for the classical Lie types. Drinfeld-Kohno theorem connects via a Drinfeld twist, the quasi-Hopf algebra $U({\mathfrak{g}})[[h]]$ with the quantized universal enveloping algebra $U_{h}({\mathfrak{g}})$ which is a Hopf algebra. As above mentioned, rigidity of Drinfeld category is derived from the equivalence constructed in Drinfeld-Kohno theorem and rigidity of the representation category of $U_{h}({\mathfrak{g}})$.
Major work was carried out in a series of papers by Kazhdan and Lusztig in the mid 90s in the setting of affine Lie algebras. The authors construct rigid braided tensor structures for certain module categories of affine Lie algebras associated to ${\mathfrak{g}}$ at levels excluding the positive rationals. In their work they consider both generic and non generic values of the levels. In particular they consider the negative integer levels $-k$, $k\in{\mathbb{N}}$, and proved a braided tensor equivalence of their category ${\mathcal{O}}_{-k}$ with the non-semisimple category of certain modules of the Drinfeld-Jimbo quantum group $U_{q}({\mathfrak{g}})$ at the same type of roots of unity $q=e^{i\pi/d\ell}$, with the exception of some exceptional cases at low levels [66]. The order $\ell$ appearing in the root of unity and the positive integer $k$ are related by a well known equality, see e.g. Def. 20.1 and 21.2 in [13] and references therein.
@@
The study of how unitarization from the quantum group fusion categories [99] was useful to derive unitarization in representation categories of vertex operator algebras via an explicit equivalence that passes through weak quasi-Hopf algebras was important for us since the early stages of applications of the theory of weak quasi-Hopf algebras [13], and further motivated us to study the above problem. Indeed one of our first results on unitarization was that weak quasi-Hopf algebras associated to fusion categories known to be equivalent allow to transport the unitary structure of one category to unitarize the other. In the process, knowledge of the explicit equivalence gives an understanding of the transported unitary structure. We have results on the structure of the unitary structure of the affine vertex operator algebra at positive integer levels transported by our equivalence, which asserts that it must correspond to the classical compact real form, and necessarily equals the unitary structure previously directly constructed by Gui, see [13] for more details.
@@
Rigidity, called charge conjugation in this setting, describes particle-antiparticle duality. In their papers rigidity is derived as a weak limit of inner automorphisms induced by transfer operators moving the support of the given localized endomorphism $\rho$ to (space-like) infinity. This gives a left inverse $\Phi_{\rho}$ of $\rho$, which is a completely positive map. From this the conjugate localized endomorphism $\overline{\rho}$ is derived. The left inverse may be described in the category as a map induced by the adjoint action of a solution of the conjugate equations on a intertwiner translated to the right by the tensor product. It was shown that the values of the statistics parameter (which is the constant $\Phi_{\rho}(\varepsilon)$, with $\varepsilon$ the action of the generator of the permutation group in the algebra of intertwining operators of the tensor powers of $\rho$), can be completely classified, and it can only take the values given by the reciprocal of a positive integer $d$ (the order of the parastatistics) up to a sign which classifies charges of para-Bose or para-Fermi type. This analysis also needed to prove that the statistics parameter can not take the value 00 and was excluded by physically motivated conditions [20], [48].
Because of the representation of the permutation group in the category, together with their results on the classification of the statistics parameter, and the $C^{*}$-structure, in the 70s it arose the question of whether that was the category of representations of a compact group, because those data arising from a single localised endomorphism were also satisfied by the representation category of a compact Lie subgroup of ${\rm SU}(d)$, that would be a gauge group of global symmetries. They considered the abelian case in the 70s [19]. In that case, every object had dimension $1$, and in modern terminology these categories are referred to as pointed in e.g. [29].
@@
Our approach may be regarded in the setting of Connes noncommutative geometry [15], and follows naturally our previous work in the setting of compact quantum groups, see also the introduction in [14], [16]. In [83], [84], [88], [85], [86], [87], [89] we have developed at the level of rigid tensor $C^{*}$-categories with simple unit an analogue of Doplicher-Roberts construction of a compact gauge group and field algebra, or more precisely the analogue of [23] in the setting of compact quantum groups. The analogy cannot be performed in a strict way due to new difficulties in the quantum case, see the mentioned papers and [13]. One important difference was that for rigid braided tensor categories we do not have braided tensor functors to the categories of vector spaces, beyond the symmetric ones, because the category of vector spaces does not admit a braided symmetry which is not a permutation symmetry. Giving up a compatibility property in the symmetric case implies lack of uniqueness of the functor, and thus of the gauge group. This was noticed early as an easy consequence from their construction.
@@
To treat the difficulty of finding a replacement of Doplicher-Roberts symmetric functors, in [12] and [13] we have taken into account Drinfeld notion of coboundary symmetry $\overline{R}$, which plays a role in the construction of Drinfeld twist in Drinfeld-Kohno theorem [27]. This is a symmetry of the category obtained modifying the braided symmetry defined by the $R$-matrix by a square root of the $2$-coboundary defined by the ribbon structure, and by construction the coboundary symmetry is an involutive symmetry, similarly to the permutation symmetry. This symmetry finally allows for a generalization of the notion of symmetric functor. Moreover we have followed both the original choice by Mack and Schomerus [71] of a given dimension function corresponding to non-negligible tensor powers of the vector representation of quantum ${\mathfrak{s}l}_{2}$ at roots of unity and also the study of classification of compact quantum groups corresponding to a given representation category of deformation compact quantum groups by Neshveyev and Yamashita [81], [82]. They both considered a specific positive integer valued dimension function associated to the given category of the quantum group, which was the classical dimension function.
@@
This suggested to us to consider in [12] and [13] a specific dimension function on a functor related to classical data, and this was first Wenzl functor [99] for the quantum group fusion category now at roots of unity, and then the minimum energy functor for the WZW model [106] in the setting of vertex operator algebras. It was particularly important for us that in [99] Wenzl describes an explicit quantization curve associated to his functor that leads to an algebra that identifies with the Zhu algebra for that model in the setting of vertex operator algebras. Since in [12] for the type $A$ case we had constructed a natural quantum group structure on a semisimple associative algebra from the quantum group fusion category endowed with Wenzl functor, compatible with the tensor structure, this arose the question of constructing another natural quantum group structure on the Zhu algebra at the time of [12], following the quantization curve on the unitary structure of tensor products with the aim of making it classical. This involved in [13] the construction of a Drinfeld twist in our setting, the above mentioned notion that Drinfeld had introduced to study equivalence of the Drinfeld category with the quantum group category in a formal setting. We have done this for all the Lie types via our version of Drinfeld-Kohno Theorem 30.4 in [13] and applications to affine vertex operator algebras at positive integer level, via Zhu linear equivalence, in [13].
@@
We refer to Sect. 25 in [13] for an extension of certain results by Neshveyev and Yamashita in the setting of quantum group fusion categories of type $A$ based on Kazhdan-Wenzl theory at roots of unity in our case, and on the weak Hopf algebras constructed in [12], in place of the compact quantum groups. This section is important for us as it suggested that in general the associativity morphisms in a category arising from CFT may be of a simple nature from the viewpoint of the associated quantum gauge groups that need to be constructed, and this finally proved to be the case, we shall come back to this with more precision.
On the other hand, in the process of extending Doplicher-Roberts theory in the setting of compact quantum groups, we had started by embedding in a tensorial way part of the morphisms of the category into the Hilbert spaces. This was the tensor subcategory generated by solutions of the conjugate equations [83], [85], [89]. Via Tannaka-Krein-Woronowicz duality this subcategory corresponds to Wang and Van Daele orthogonal or universal compact quantum groups $A_{o}(F)$, or $A_{u}(F)$ [97], their representation category has been studied by Banica [2], [3].
@@
The choice of the compact quantum group arising in this way depends on whether a generating object $\rho$ is selfconjugate as in the case of $A_{o}(F)$ or not as for $A_{u}(F)$. In doing this embedding we had already made a choice of a dimension function, the rank of the associated invertible matrix $F$, and from that choice we had constructed a tensor functor from that subcategory. To consider the full category, the analogue Doplicher-Roberts compact gauge group is described by an ergodic action of the corresponding quantum group on a quantum homogeneous space rather than a quantum group itself, due to lack of commutativity of this $C^{*}$-algebra [85]. The homogeneous space is described by a $C^{*}$-algebra, naturally associated to the category and to the embedding of the subcategory.
This property is likely to hold also in other models arising from conformal field theory starting from the minimum energy functor, unless one can show that certain known obstructions already met in [84] for compact quantum groups, vanish, we shall come back to this in the outlook section 13.
@@
However, for the WZW models arising from affine Lie algebras at positive integer levels in [12] for the type $A$ case, and in [13] for all the Lie types, we obtain genuine quantum groups as quantum gauge groups, described as semisimple weak Hopf algebras with $R$-matrix, $C^{*}$-structure explicitly associated to Drinfeld coboundary matrix, and with unitary representations of the braid group, derived from the quantum group fusion category. We prove that these weak Hopf algebras have representation category equivalent to both the starting fusion category for all the Lie types and also to that of the corresponding affine vertex operator algebra with ribbon braided Huang-Lepowsky tensor structure, for the classical Lie types and $G_{2}$. Constructing the equivalence from the former to the latter in the setting of semisimple weak quasi-Hopf algebras as we do in [13] for these Lie types is our proof of Kazhdan-Lusztig-Finkelberg theorem, see Theorem 2.1.
@@
Among our first references in the context of conformal field theory important for us to start to set up the connection with the work of Wenzl in quantum groups, is the work by Wassermann [98] in the setting of affine Lie algebras, where the key concept is played by the notion of primary field, and its initial term, which we use in a fundamental way to connect quantum group fusion category to representations of the Zhu algebra and their special tensor products with the fundamental representation, see Sect. 35 in [13]. Most importantly, we describe a connection between Wenzl and Wassermann work via Wenzl quantisation curve, and building on their work in [13] we obtain Theorems 3.1, 3.2, Cor. 3.3, and Theorem 3.4 which compares with the tensor product structure by Huang and Lepowsky, to which we refer the reader also for important historical references. The connection from affine Lie algebra setting to Huang-Lepowsky tensor product theory may be found in Sections 34–39 in [13].
@@
Our abstract duality methods need new further tools as we shall see from the theory of quantum groups, mainly from the work of Drinfeld and Wenzl, that are not available in the theory of groups. However, we shall find that in both contexts, quantum groups at roots of unity and affine vertex operator algebras at positive integer levels, both representation categories are described by functors extending the symmetric tensor functors in Doplicher-Roberts abstract duality theory, see Sect. 29 in [13]. We call these structures hermitian coboundary weak quasi-Hopf algebras, as the hermitian structure is given by Drinfeld coboundary matrix, which is a modification of the $R$-matrix with a square root of the ribbon structure.
@@
We notice on the other hand that a stronger property is needed to establish the connection, than the purely categorical characterisation of functors in Sect. 29 in [13]. This property is expressed in terms of quantum groups, and the first example, in a non-semisimple setting, is the hermitian structure of $U_{q}({\mathfrak{g}})$ at a root of unity by Theorem 31.9 in [13]. This property says that the -involution enjoys a precise anticommutation relation (rather than commutation relation as is familiar in Hopf $C^{*}$-algebras) with the coproduct and was first noticed in [99]. We axiomatize this property in Sect. 30 in [13], as a compatibility property of hermitian coboundary weak quasi-Hopf algebras. Further examples are our weak bi-algebras $A_{W}({\mathfrak{g}},q)$ associated to Wenzl functor $W:{\mathcal{C}}({\mathfrak{g}},q)\to{\rm Vec}$ and to the quasi-bialgebras structures on the Zhu algebra $A({V_{{\mathfrak{g}}_{k}}})$ associated to Zhu functor ${\rm Rep}({V_{{\mathfrak{g}}_{k}}})\to{\rm Vec}$ via our Drinfeld twist construction, see Sections 32, 34 in [13]. The anticommutation property of the -involution and coproduct of $A_{W}({\mathfrak{g}},q)$ is not a direct consequence of the corresponding property of $U_{q}({\mathfrak{g}})$, as its coproduct is not compatible with the fusion tensor product. We need some delicate analysis of the fusion tensor product described in [99]. For $A({V_{{\mathfrak{g}}_{k}}})$, we show that the twist makes the coproduct of $A({V_{{\mathfrak{g}}_{k}}})$ symmetric, and this is a result analogous to the symmetry property of the usual coproduct of Drinfeld quasi-Hopf algebra $U({\mathfrak{g}})[[h]]$. For details we refer to the cited sections.
On the other hand, as an algebra $A_{W}({\mathfrak{g}},q)$ is a natural quotient of $U_{q}({\mathfrak{g}})$, and the antipode of $A_{W}({\mathfrak{g}},q)$ is directly derived from that of $U_{q}({\mathfrak{g}})$. In this way, $A_{W}({\mathfrak{g}},q)$ becomes a weak Hopf algebra. It follows that ${\rm Rep}(A_{W}({\mathfrak{g}},q))$ becomes a rigid ribbon braided tensor category naturally equivalent to ${\rm Rep}(A({V_{{\mathfrak{g}}_{k}}}))$ via our twist, and therefore so does ${\rm Rep}({V_{{\mathfrak{g}}_{k}}})$ via Zhu equivalence.
@@
As above motivated, we shall derive more structure beyond rigidity and modularity on ${\rm Rep}({V_{{\mathfrak{g}}_{k}}})$ from the quantum group fusion category, a further example is unitarization of the vertex operator algebra representation category from the corresponding quantum group fusion category, via the explicit equivalence, see Sects. 22, 23, 24, 26 in [13]. To this aim it is important that our equivalence passes through the weak quasi-Hopf algebras, and the fact that these are naturally associated to the categories. Indeed it was known that a mere tensor equivalence between the categories would not suffice to transport the unitary structures see [49]. Sect. 16 in [13] treats this for general tensor categories that admit submultiplicative dimension functions (called weak dimension functions).
@@
Our equivalence, since it is given by a Drinfeld twist analogous to [25], [27], and is also motivated by the requirement of making the unitary structure classical, at the level of the associated weak Hopf or quasi-Hopf algebras gives an understanding of the unitary structures of representation categories of affine vertex operator algebras as classical compact real forms, or untwisted. We refer to the outlook section for some more information on possible extension of our viewpoint to models beyond the affine vertex operator algebras at positive integer levels, see also [13].
@@
Our main result is Theorem 2.1. In the subsequent sections we outline the plan of our proof. A main idea that we follow is to consider the fundamental representation $V$ of ${\mathfrak{g}}$ for each Lie type as in [99] where the author explains its generating property and good fusion structure with respect to tensor products with an arbitrary irreducible representation allowing a well-defined truncation procedure to obtain non-negligible objects, with the aim of studying unitarizability of the fusion category. It should also be noted that a main application of that paper is construction of inclusion of subfactors of finite Jones index [63] for all Lie types, extending his previous work.
@@
For all the Lie types we obtain complete identification for the tensor product structure, all the ribbon structure, and those braiding and associativity morphisms where one or two variables are the fundamental representation $V$, respectively, with Huang-Lepowsky ribbon braided tensor structure. These are parts (a), (b), (c) of Theorem 2.1.
@@
In [12], [13], and also in this paper, weak Hopf algebras are understood with a different meaning. While a different name would be helpful to avoid confusion, so far we have not been able to find a good name for our algebras to clarify the situation, and we use the term weak Hopf algebras in the sense defined in Sect. 10 in [13].